Consider the system: I’ – 41+y” = ť? I+I+y = 0 In the first blank: Write the second equation using the D notation. (Do not
Answer #1

The given system is, x_4x of y = ² xtaty=0 (1 72) Here at at in Do notation is so the 2nd eq equation Datap Dy=0 (D+1)x ofNow particular integral, aplt) = (-6) D² + 4 () 1(1+B) -- (142) Sky =-11-+)() -#144.2) --$145) - #(4+) X(t) = 4 652+ + ca sin